# SoK: Large Language Model Copyright Auditing via Fingerprinting

### v2

@@ 3.1 Framework of LLM Fingerprinting
The framework hinges on the careful design of these two functions. The $\text{Extract}(\cdot)$ function needs to isolate a signal that is sensitive to a model’s provenance. The $\text{Verify}(\cdot)$ function then measures this signal to make a final, decisive judgment. Together, they provide a systematic methodology for navigating the complex challenge of LLM copyright auditing.
@@ 3.2 Design Objectives
  • Discriminability: To achieve both high effectiveness and uniqueness, the fingerprint should produce distributions for derivative and independent models that are maximally separated in the feature space, enabling reliable differentiation.
@@ 3.3 Taxonomy of LLM Fingerprinting
In this paper, we classify LLM fingerprinting methods based on a single, practical criterion: the level of access an auditor has to the suspicious model, $M_{s}$. This distinction fundamentally dictates the techniques that can be applied. Existing LLM fingerprinting methods fall into one of two categories: white-box [76, 79] or black-box fingerprinting [44, 26].
  • White-box LLM Fingerprinting: This category assumes the auditor has full access to the suspicious model’s internal components. This includes its architecture and parameters. With this level of access, fingerprints can be extracted by directly analyzing the model’s static weights or its dynamic internal states during forward or backward passes.
@@
This distinction in access level fundamentally shapes the practical applicability of each category. White-box fingerprinting methods are best suited for auditing open-source models, where auditors can directly download and examine model parameters. In contrast, black-box fingerprinting methods apply to a broader set of scenarios, including the auditing of proprietary commercial models accessed via APIs, as well as cases of model theft in which an adversary deploys a stolen model as a closed service. A summary of existing approaches is provided in Appendix A.
@@
4 White-box LLM Fingerprinting
White-box LLM fingerprinting methods operate under the assumption of full access to a model’s internal architecture and parameters. We classify these methods into three different categories, static, forward-pass, and backward-pass fingerprinting. This taxonomy is determined by whether the features are derived from the model in a static state, during a forward pass, or during a backward pass, as shown in Figure 1.
4.1 Static Fingerprinting
Static fingerprinting methods are the most direct form of white-box auditing. They operate directly on the model’s learned parameters $\{W_{1},\dots,W_{k}\}$, treating the static weights themselves as the primary source of identifying information.
Architecture-agnostic Fingerprinting. Some static fingerprinting approaches treat a model’s parameters as a single undifferentiated tensor of numerical values and extract fingerprints from all or part of the model weights. This strategy is architecture-agnostic and does not depend on any specific model design. For example, a simple baseline method, WeightCompare[34], measures model similarity by computing a score based on the number of layers with identical structures and weights across two models. Beyond such direct comparisons, more advanced techniques aim to generate fingerprints that are both robust and efficient. Zheng et al.[85] proposed one such method, which produces a non-repudiable fingerprint by projecting front-layer weights onto a random space defined by the model owner’s unique identity. This yields a compact and irrevocable fingerprint capable of identifying pirated models, even those retrained for different applications.
Transformer-specific Fingerprinting. Currently, modern LLMs are mostly based on the Transformer architecture [61]. Based on this understanding, a more targeted method tailors the fingerprinting process to this specific structure. Instead of viewing parameters as a generic block of data, these methods selectively extract features from key architectural components ($i.e.$, query, key, value, and output matrices) of transformers. Zeng et al. [76] proposed HuRef, which constructs invariant terms by combining weight matrices from attention and feed-forward layers, designing them to be resilient against manipulations like parameter permutation. More straightforwardly, Yoon et al. [75] discover that the standard deviation distributions of attention matrices across layers form a stable intrinsic fingerprint robust to continued training. Recently, Zhang proposed MDIR [81], a method using matrix analysis to detect and reconstruct weight relationships. MDIR leverages polar decomposition to compare the orthogonal parts of weight matrices and can identify the exact transformation between them, providing a statistical p-value for the detected similarity using Large Deviation Theory.
@@ 4.2 Forward-pass Fingerprinting
Forward-pass fingerprinting methods [7, 78, 79] leverage the dynamic internal states of an LLM as it processes input data. The fundamental idea is that during a forward pass, a model generates a vast number of intermediate representations $\{R_{1},\dots,R_{k}\}$, which are shaped by its unique architecture and learned parameters. These rich, high-dimensional features can serve as a distinctive fingerprint to characterize the model’s behavior and origins. Unlike static methods that analyze model weights in isolation, forward-pass techniques capture the functional properties of a model in action.
Recent works have proposed various instantiations of forward-pass fingerprinting. A pioneering work, DeepJudge by Chen et al. [7], introduced a comprehensive testing framework using multiple metrics on intermediate model outputs. It generates specialized test cases, such as adversarial examples or inputs that could trigger corner-case neuron activations, to amplify differences, which are then quantified using direct distance measures. Following DeepJudge, EasyDetector [78] employs an indirect comparison method by training a simple linear classifier (also called “linear probe”) on the source model’s representations; the probe’s classification accuracy on a suspect model then serves as the similarity score. To enhance robustness, REEF [79] uses Centered Kernel Alignment (CKA) to compare the geometry of representation spaces, creating a fingerprint resilient to common modifications like model pruning and parameter permutation.
@@ 4.3 Backward-pass Fingerprinting
To the best of our knowledge, the only existing work on backward-pass fingerprinting for LLMs is TensorGuard [62]. Its core idea is to characterize a model using statistical features of gradients ($e.g.$, mean, standard deviation, norm). To generate a stable and robust signature, TensorGuard perturbs the model by injecting controlled random noise into internal tensors during a forward pass, which induces consistent gradient responses captured via backpropagation. Statistical features are then extracted from these gradients, and principal component analysis (PCA) is applied for dimensionality reduction to enhance efficiency. Finally, similarity is measured by computing the Euclidean distance between the resulting fingerprint vectors. This method yields a stable signature that captures the model’s intrinsic gradient-based behaviors.
@@ 5 Black-box LLM Fingerprinting
Black-box LLM fingerprinting methods operate under the constraint that the auditor has no access to the suspect model’s internal components, such as its architecture or parameters. Instead, the auditor can only interact with the model through its public API by issuing queries and analyzing the corresponding responses. The central challenge in this setting is to derive a unique and reliable fingerprint solely from observable input–output behavior. In our taxonomy, we further classify black-box fingerprinting into two categories: untargeted and targeted, depending on whether the method relies on developer-predefined input–output pairs.
@@ 5.1 Untargeted Fingerprinting
Untargeted fingerprinting [72, 24, 44] represents a direct and intuitive approach to black-box LLM copyright auditing. Several preliminary studies [53, 3, 55] observed that the outputs generated by different LLMs are often distinguishable, exhibiting unique stylistic or behavioral “idiosyncrasies”. Inspired by this observation, untargeted fingerprinting methods operate by first selecting a set of queries and taking the (features of) corresponding responses as fingerprints. A high degree of similarity between the responses of different LLMs indicates that one is a derivative of the other.
@@
Choice of the Query Set $Q$. The choice of the query set $Q$ is critical as it determines the nature of the behavioral signals elicited from the model. Existing works have explored several distinct strategies for selecting or generating these queries.
  • Manually Constructed Query Sets: This method uses queries that are either strategically handcrafted or sampled from existing datasets. Methods like LLMmap [44] and DuFFin [70] manually curate queries that target specific model behaviors, such as alignment and knowledge across various domains. To ensure stealth, methods like model equality testing (MET) [13] and Rank-Based Uniformity Test (RUT) [86] instead use prompts from general corpora like Wikipedia or user-chatbot conversations to mimic natural user interaction patterns.
  • Automatically Constructed Query Sets: To discover more discriminative queries, some works also explore automating the generation process. For instance, Hide-and-Seek [24] employs an evolutionary algorithm where an LLM auditor iteratively generates and refines prompts to maximize response differences among various models. Alternatively, Kurian et al. [30] employ reinforcement learning to automatically select the optimal query subset from a larger, pre-defined pool.
  • Query Set Exploring Specific LLM Capabilities: A third method focuses on fingerprinting a single, complex capability of an LLM. For instance, CoTSRF [47] uses a query set of reasoning questions paired with a Chain-of-Thought (CoT) trigger to specifically probe and capture the model’s unique logical reasoning patterns.
@@
  • Using Output Logits or Probabilities: The most direct methods operate on a model’s output logits or probabilities, such as reconstructing a unique vector space from the probability distribution [72] or computing a scalar log-rank score for each response [86].
  • Leveraging Lexical and Syntactic Features: Other methods analyze the generated text for its lexical and syntactic features [2]. For instance, McGovern et al. [42] show that character, word, and part-of-speech (POS) n-grams could serve as a robust fingerprint.
  • Employing Neural Networks: The most prevalent techniques employ neural networks to learn a feature embedding. Methods including LLMmap [44], CoTSRF [47], DuFFin [70], and Kurian et al.’s method [30] all leverage Transformer-based models to encode responses into a dense fingerprint vector that captures the model’s unique features.
Design of the Similarity Function $\rm{Sim}(\cdot,\cdot)$. Finally, the similarity function $\rm{Sim}(\cdot,\cdot)$ compares the extracted fingerprints to decide model identity, with existing methods ranging from simple distance measures to formal statistical tests. The most straightforward approaches compute a distance or similarity score directly in the feature space, using metrics such as cosine similarity [44, 70], Hamming distance [70], or distributional KL-divergence [47]. More advanced approaches formulate the task as a classification problem, training a model—such as a GradientBoost classifier [42] or a dedicated Transformer network [44, 2]—to map fingerprints back to their source LLMs. At the most formal level, fingerprint comparison is treated as a statistical two-sample hypothesis test. For example, Gao et al.[13] employ a Maximum Mean Discrepancy (MMD) test, while RUT[86] applies a Cramér–von Mises test to rigorously determine whether two sets of fingerprints are drawn from the same distribution.
@@ 5.2 Targeted Fingerprinting
In contrast to untargeted fingerprinting, targeted fingerprinting seeks to construct a specific set of query–response pairs that are unique to a source model and its derivatives [17, 26, 66, 60]. The key idea is that these pairs, collectively forming the fingerprint, will be consistently reproduced only by models in the source lineage, whereas independently trained models will fail to generate the corresponding target responses. The essential distinction between untargeted and targeted approaches lies in whether fingerprint verification relies on predefined query–response pairs. Notably, targeted methods yield an asymmetric fingerprint: for a given targeted method $\mathcal{A}$, the outcome of $\mathcal{A}(M_{o},M_{s})$ is not necessarily equal to $\mathcal{A}(M_{s},M_{o})$.
@@
The primary innovation within targeted fingerprinting lies in methods for generating unique query-response pairs, often using gradient-based optimization inspired by adversarial attacks [87]. Methods like TRAP [17] and ProFLingo [26] craft an adversarial prompt component to force a source model to produce a specific, pre-determined response, such as a designated number or a factually incorrect answer. To enhance the robustness of this fingerprint across an entire model lineage, RAP-SM [66] extends this technique by using shadow models ($i.e.$, known derivatives of the source model) for joint optimization. Furthermore, a distinct method, RoFL [60], generates fingerprints without a pre-defined target response. Instead, it first gets a response by feeding an unlikely token sequence to the source model. Then RoFL optimizes the prompt to consistently re-elicit the same response across model variants.
@@ 6 LeaFBench: Benchmarking LLM Fingerprinting in the Wild
While numerous LLM fingerprinting techniques have been proposed, they are often evaluated under disparate settings, making direct and fair comparison a significant challenge. To date, the field has still lacked a unified and standardized benchmark for assessing the reliability and robustness of these methods. To address this gap, this paper introduces LeaFBench, an initial benchmark designed to facilitate a more systematic evaluation of LLM fingerprinting techniques against a curated set of realistic model modifications and deployment scenarios. Besides, LeaFBench provides extensible and user-friendly interfaces, enabling the evaluation of new fingerprinting methods and additional post-development techniques.
@@ 6.1 Techniques affecting LLM Fingerprinting
After an LLM is pre-trained, it often undergoes a variety of post-development modifications before being deployed in a real-world application. These techniques, designed to adapt or optimize the model, can significantly alter its behavior. Consequently, the stability and detectability of its fingerprint may be impacted. We broadly classify these influential techniques into two categories: parameter-altering techniques and parameter-independent techniques.
@@
  • Fine-tuning. It is the process of further training a pre-trained model on a smaller and task-specific dataset ($i.e.$, personalization dataset) to specialize its capabilities [22].
  • Quantization. It is a compression technique that reduces the precision of model weights ($e.g.$, from 32 bits to 8 bits), thereby making the model smaller and faster [28].
  • Distillation. It is used to train a smaller “student” model for mimicing a larger “teacher” model, transferring knowledge into a more compact form [71].
  • Model Merge. It is a technique that combines the weights of two or more fine-tuned models to create a new model inheriting capabilities from its predecessors [15].
@@
  • System Prompts: System prompts are high-level instructions provided to the model alongside the user’s query to set a specific context or set of rules for the interaction [14].
@@
  • Sampling Strategies: Sampling strategies are algorithms that control how the next token is selected from the model’s output probability distribution. Common methods include Temperature scaling, Top-k, and Top-p sampling, which adjust the randomness and diversity of the generated text [77].
@@
  • Retrieval-Augmented Generation (RAG): RAG enhances model responses by retrieving relevant information from an external knowledge base and supplying it as additional context during generation [31].
6.2 Overview of LeaFBench

### v3

@@ 3.1 LLM Fingerprinting Framework
The framework hinges on the careful design of these two functions. The $\text{Extract}(\cdot)$ function needs to isolate a signal that is sensitive to a model’s provenance. The $\text{Verify}(\cdot)$ function then measures this signal to make a final, decisive judgment. Together, they provide a systematic methodology for navigating the challenge of LLM copyright auditing.
@@ 3.2 Design Objectives
  • Discriminability: To achieve both high effectiveness and uniqueness, the difference between fingerprints of derivative and independent models should be maximized, enabling reliable differentiation.
@@ 3.3 Taxonomy of LLM Fingerprinting
In this paper, we first classify existing LLM fingerprinting methods based on a practical criterion: the level of access an auditor has to the suspicious model, $M_{s}$. This primary distinction is crucial because the information accessible in white-box and black-box scenarios differs fundamentally, leading to significant disparities in their technical paradigms and evaluation methods. Based on this, existing techniques fall into one of two categories: white-box [zeng2024huref, zhang2025reef] or black-box LLM fingerprinting [pasquini2025llmmap, jin2024proflingo].
  • White-box LLM Fingerprinting: This category assumes the auditor has full access to the suspicious model’s internal components, including its architecture and parameters. With this level of access, fingerprints can be extracted by directly analyzing the model’s static weights or its dynamic internal states.
@@
This distinction in access level fundamentally shapes the practical applicability of each category. White-box fingerprinting methods are best suited for auditing open-source models, where auditors can directly download and examine model parameters. In contrast, black-box fingerprinting methods apply to a broader set of scenarios, including the auditing of proprietary commercial models accessed via APIs, as well as cases of model theft in which an adversary deploys a stolen model as a closed service.
Building upon this foundational division, we will further provide a detailed sub-classification for both white-box and black-box approaches. We classify white-box methods into three different subcategories, static, forward-pass, and backward-pass fingerprinting, and black-box methods into two different subcategories, untargeted and targeted fingerprinting. Detailed introductions are in Section 4&5. A summary of existing approaches is provided in Appendix .1.
4 White-box LLM Fingerprinting
White-box LLM fingerprinting methods operate under the assumption of full access to a model’s internal architecture and parameters. We classify these methods into three different categories, static, forward-pass, and backward-pass fingerprinting. This taxonomy is determined by whether the features are derived from the model in a static state, during a forward pass, or during a backward pass, as Figure 1.
4.1 Static Fingerprinting
Static fingerprinting methods [li2021modeldiff, zheng2022dnn, zeng2024huref, yoon2025intrinsic] are the most straightforward form of white-box auditing. They operate directly on the model’s learned parameters $\{W_{1},\dots,W_{k}\}$, treating the static weights themselves as the primary source of identifying information.
Architecture-agnostic Fingerprinting. Some static fingerprinting approaches treat a model’s parameters as a single undifferentiated tensor of numerical values and extract fingerprints from all or part of the model weights. A simple method, WeightCompare [li2021modeldiff], measures model similarity by computing a score based on the number of layers with identical structures and weights across two models. Beyond such direct comparisons, the more advanced technique aims to generate fingerprints that are more robust and efficient. Zheng et al. [zheng2022dnn] proposed one such method, which produces a non-repudiable fingerprint by projecting front-layer weights onto a random space defined by the model owner’s unique identity. These methods are architecture-agnostic and do not depend on any specific model design.
@@ 4.1 Static Fingerprinting
Transformer-specific Fingerprinting. Currently, modern LLMs are mostly based on the Transformer architecture [vaswani2017attention]. Based on this understanding, a more targeted method tailors the fingerprinting process to this specific structure. Instead of viewing parameters as a generic block of data, these methods selectively extract features from key architectural components ($i.e.$, query, key, value, and output matrices) of transformers. Zeng et al. [zeng2024huref] proposed HuRef, which constructs invariant terms by combining weight matrices from attention and feed-forward layers, designing them to be resilient against manipulations like parameter permutation. More straightforwardly, Yoon et al. [yoon2025intrinsic] discover that the standard deviation distributions of attention matrices across layers could form an intrinsic fingerprint. Recently, Zhang proposed MDIR [zhang2025matrix], a method using matrix analysis to detect and reconstruct weight relationships. MDIR leverages polar decomposition to compare the orthogonal parts of weight matrices and can identify the exact transformation between them, providing a statistical p-value for the detected similarity using Large Deviation Theory.
@@ 4.2 Forward-pass Fingerprinting
Forward-pass fingerprinting methods [chen2022copy, zhang2024easydetector, zhang2025reef] leverage the dynamic internal states of an LLM as it processes input data. The fundamental idea is that during a forward pass, a model generates a vast number of intermediate representations $\{R_{1},\dots,R_{k}\}$ ($k$ is the number of layers), which are shaped by its unique architecture and learned parameters. These high-dimensional features can serve as a distinctive fingerprint to characterize the model’s behavior and origins. Unlike static methods that analyze model weights in isolation, forward-pass techniques capture the functional properties of a model in action.
Recent works have proposed various instantiations of forward-pass fingerprinting. A pioneering work named DeepJudge [chen2022copy] introduced a comprehensive testing framework using multiple distance metrics on intermediate model outputs. It generates specialized test cases, such as adversarial examples or inputs that could trigger corner-case neuron activations, to amplify differences, which are then quantified using direct distance measures. Following DeepJudge, EasyDetector [zhang2024easydetector] employs an indirect comparison method by training a simple linear classifier (also called “linear probe”) on the source model’s representations. The probe’s classification confidence on a suspect model then serves as the similarity score. To enhance robustness, REEF [zhang2025reef] uses Centered Kernel Alignment to compare the geometry of different models’ representations.
@@ 4.3 Backward-pass Fingerprinting
To the best of our knowledge, the only existing work on backward-pass fingerprinting for LLMs is TensorGuard [wu2025gradient]. Its core idea is to characterize a model using statistical features of gradients ($e.g.$, mean, standard deviation, norm). To generate a stable and robust signature, TensorGuard perturbs the model by injecting controlled random noise into internal tensors during a forward pass, which induces consistent gradient responses captured via backpropagation. Statistical features are then extracted from these gradients, and principal component analysis is applied for dimensionality reduction to enhance efficiency. Finally, similarity is measured by computing the Euclidean distance between the resulting fingerprint vectors.
@@ 5 Black-box LLM Fingerprinting
Black-box LLM fingerprinting methods operate under the constraint that the auditor has no access to the suspect model’s internal components, such as its architecture or parameters. Instead, the auditor can only interact with the model through its public API by issuing queries and analyzing the corresponding responses. The central challenge in this setting is to derive a unique and reliable fingerprint solely from observable input–output behavior. In our taxonomy, we further classify black-box fingerprinting into two categories: untargeted and targeted, depending on whether the method relies on detecting developer-predefined input–output pairs.
@@ 5.1 Untargeted Fingerprinting
Untargeted fingerprinting [yang2024fingerprint, iourovitski2024hide, pasquini2025llmmap] represents a direct and intuitive approach to black-box LLM copyright auditing. Several preliminary studies observed that the outputs generated by different LLMs are often distinguishable, exhibiting unique stylistic or behavioral “idiosyncrasies” [sun2025idiosyncrasies, bitton2025detecting, suzuki2025natural]. Inspired by this observation, untargeted fingerprinting typically operates by first selecting a set of input queries and taking the (features of) corresponding output responses as fingerprints. A high degree of similarity between the responses of different LLMs could indicate that one is a derivative of the other.
@@
Choice of the Query Set $Q$. The choice of the query set $Q$ is critical as it determines the nature of the behavioral signals elicited from the models. Existing works have explored several strategies for selecting or generating these queries:
  • Manually Constructed Query Sets: This method uses queries that are either strategically handcrafted or sampled from existing datasets. Methods like LLMmap [pasquini2025llmmap] and DuFFin [yan2025duffin] manually curate queries that target specific model behaviors, such as alignment and knowledge across various domains. Other methods like model equality testing (MET) [gao2025model] and Rank-Based Uniformity Test (RUT) [zhu2025auditing] instead use prompts from general corpora like Wikipedia or user-chatbot conversations to mimic natural user interaction patterns.
  • Automatically Generated Query Sets: To discover more discriminative queries, some works also explore automating the generation process. For instance, Hide-and-Seek [iourovitski2024hide] employs an evolutionary algorithm where an LLM iteratively generates and refines prompts to maximize response differences among various models. Alternatively, Kurian et al. [kurian2025attacks] employ reinforcement learning to automatically select the optimal query subset from a larger, pre-defined pool.
  • Query Set Exploring Specific LLM Capabilities: A third method focuses on fingerprinting a single, complex capability of an LLM. For instance, CoTSRF [ren2025cotsrf] uses a query set of reasoning questions paired with a Chain-of-Thought (CoT) trigger to specifically probe and capture the model’s unique logical reasoning patterns.
@@
  • Using Output Logits or Probabilities: The most direct methods operate on a model’s output logits or probabilities, such as reconstructing a unique vector space from the probability distribution [yang2024fingerprint] or computing a scalar log-rank score for each response [zhu2025auditing].
  • Leveraging Lexical and Syntactic Features: Other methods analyze the generated text for its lexical and syntactic features [bhardwaj2025invisible]. For instance, McGovern et al. [mcgovern2025your] show that character, word, and part-of-speech (POS) n-grams could serve as a robust fingerprint.
  • Employing Neural Networks: The most prevalent techniques employ neural networks to learn a feature embedding. Methods including LLMmap [pasquini2025llmmap], CoTSRF [ren2025cotsrf], DuFFin [yan2025duffin], and Kurian et al.’s method [kurian2025attacks] all leverage Transformer-based models to encode responses into a dense fingerprint vector that captures the model’s features.
Design of the Similarity Function $\rm{Sim}(\cdot,\cdot)$. Finally, the similarity function $\rm{Sim}(\cdot,\cdot)$ compares the extracted fingerprints to decide model identity, with existing methods ranging from simple distance measures to formal statistical tests. The most straightforward approaches compute a distance or similarity score directly in the feature space, using metrics such as cosine similarity [pasquini2025llmmap, yan2025duffin], Hamming distance [yan2025duffin], or distributional KL-divergence [ren2025cotsrf]. More advanced approaches formulate the task as a classification problem, training a model ($e.g.$, a GradientBoost classifier [mcgovern2025your] or a dedicated Transformer network [pasquini2025llmmap, bhardwaj2025invisible]) to map fingerprints back to their source LLMs. At the most formal level, fingerprint comparison is treated as a statistical two-sample hypothesis test. For example, Gao et al.[gao2025model] employ a Maximum Mean Discrepancy (MMD) test, while RUT[zhu2025auditing] applies a Cramér–von Mises test to rigorously determine whether two sets of fingerprints are drawn from the same distribution.
@@ 5.2 Targeted Fingerprinting
In contrast to untargeted fingerprinting, targeted fingerprinting seeks to construct a specific set of query–response pairs that are unique to a source model and its derivatives [gubri2024trap, jin2024proflingo, xu2025rap, tsai2025rofl]. The key idea is that these pairs, collectively forming the fingerprint, will be consistently reproduced only by models in the source lineage, whereas independently trained models will fail to generate the corresponding target responses. The essential distinction between untargeted and targeted approaches lies in whether fingerprint verification relies on predefined query–response pairs. Notably, targeted methods yield an asymmetric fingerprint: for a given targeted method $\mathcal{A}$, the outcome of $\mathcal{A}(M_{o},M_{s})$ is not necessarily equal to $\mathcal{A}(M_{s},M_{o})$.
@@
The primary innovation within targeted fingerprinting lies in methods for generating unique query-response pairs, often using gradient-based optimization inspired by adversarial attacks [zou2023universal]. Methods like TRAP [gubri2024trap] and ProFLingo [jin2024proflingo] craft an adversarial prompt component to force a source model to produce a specific, pre-determined response, such as a designated number or a factually incorrect answer. To enhance the robustness of this fingerprint across an entire model lineage, RAP-SM [xu2025rap] extends this technique by using shadow models ($i.e.$, known derivatives of the source model) for joint optimization. Furthermore, a distinct method, RoFL [tsai2025rofl], generates fingerprints without a pre-defined target response. Instead, it first gets a response by feeding an unlikely token sequence to the source model. Then RoFL optimizes the prompt to consistently re-elicit the same response across model variants.
@@ 6 LeaFBench: Benchmarking LLM Fingerprinting in the Wild
While numerous LLM fingerprinting techniques have been proposed, they are often evaluated under disparate settings, making direct and fair comparison a significant challenge. To date, the field has still lacked a unified and standardized benchmark for assessing the effectiveness and reliability of these methods. To address this gap, we introduce LeaFBench, a benchmark designed to facilitate a systematic evaluation of LLM fingerprinting techniques against a curated representative set of realistic model modifications and deployment scenarios. Besides, LeaFBench provides extensible and user-friendly interfaces, enabling the evaluation of new fingerprinting methods and additional post-development techniques.
@@ 6.1 Techniques affecting LLM Fingerprinting
After an LLM is pre-trained, it often undergoes a variety of post-development modifications before being deployed in a real-world application. These techniques, designed to adapt or optimize the model, can significantly alter its parameters and behavior. Consequently, the stability and detectability of its fingerprint may be impacted. We broadly classify these influential techniques into two categories: parameter-altering and parameter-independent techniques.
@@
  • Fine-tuning. It is the process of further training a pre-trained model on a smaller and task-specific dataset ($i.e.$, personalization dataset) to specialize its capabilities [hu2022lora].
  • Quantization. It is a compression technique that reduces the precision of model weights ($e.g.$, from 32 bits to 8 bits), thereby making the model smaller and faster [jin2024comprehensive].
  • Distillation. It is used to train a smaller “student” model for mimicing a larger “teacher” model, transferring knowledge into a more compact form [yang2024survey].
  • Model Merge. It is a technique that combines the weights of two or more different models to create a new model inheriting capabilities from its predecessors [goddard2024arcee].
@@
  • System Prompts: System prompts are high-level instructions provided to the model alongside the user’s query to set a specific context or set of rules for the interaction [giray2023prompt].
@@
  • Sampling Strategies: Sampling strategies are algorithms that control how the next token is selected from the model’s output probability distribution. Common techniques include Temperature scaling, Top-k, and Top-p sampling, which adjust the randomness and diversity of the generated text [zhang2023survey].
@@
  • Retrieval-Augmented Generation (RAG): RAG enhances model responses by retrieving relevant information from an external knowledge base and supplying it as additional context during generation [lewis2020retrieval].
6.2 Construction of LeaFBench